OpenStudy (anonymous):

Calculate the length of AD given AB = 4 cm, AC =10 cm, and AD = 5. Find AE. become fan and give award

OpenStudy (jdoe0001):

is there a picture you'd like to share with us?

OpenStudy (anonymous):

yeah

OpenStudy (jdoe0001):

ok, we're anxiously waiting your excellent quality picture :)

OpenStudy (anonymous):

here

OpenStudy (anonymous):

the one on top

OpenStudy (jdoe0001):

AB * AC = AD * AE solve for AE

OpenStudy (jdoe0001):

btw the wording is a bit off "Calculate the length of \(\bf AD\) given AB = 4 cm, AC =10 cm, and \(\bf AD = 5\)" hehe

OpenStudy (anonymous):

6

OpenStudy (jdoe0001):

4 * 10 = 5 * AE

OpenStudy (jdoe0001):

is not 6 btw

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

2.5

OpenStudy (jdoe0001):

let's take a peek at the picture, hmm, AC looks awfully long, AC = 10, AE has to be almost as long, 2.5? doesn't look like

OpenStudy (anonymous):

how about 25

OpenStudy (jdoe0001):

25 is more than twice as much as 10, in the picture it doesn't look like so how about you solve by AE? :) 4 * 10 = 5 * AE

OpenStudy (anonymous):

9

OpenStudy (jdoe0001):

how did you get 9?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

my mistake

Directrix (directrix):

@gravegokuss5 Review the theorem for this problem.

Directrix (directrix):

4 * 10 = 5 * AE 40 = 5* (AE) 40/5 = AE ? = AE @gravegokuss5 Hint: AE is not 5